If 64 is added to a number, the result is 42 less than three times the number. what is the number?

Answer by checkley79(3341)

N+64=3N-42
N-3N=-42-64
-2N=-106
N=-106/-2
N=53 ANS.
PROOF:
53+64=3*53-42
117=159-42
117=117